AdvTech Limited is a large firm listed on the Johannesburg Stock Exchange. It has the following capital structure: Convertible Debt – 5 yrs; 8% equals R25 million Preferred Shares – 5% coupon + nominal value of 100 equals R15 million Common Equity (nominal value R 10/share) equals R10 million Retained Earnings R23 million The current dividend for the company is R 50/share and is expected to grow at 3% per year in the foreseeable future. The equity shares trade at R450/share. The preferred shares trade at R 104/share. The convertible debt has a conversion privilege of 2 shares per R1,000 face value at maturity. The debt currently trades at R 950. The firm’s income tax rate is 30% Required: 3.1 Calculate the firm’s weighted average cost of capital (WACC). (13 marks) 3.2 Discuss the firm’s dividend payout policy and whether it has an impact on share price. (6 marks)
Answer #1
Cost of preferred shares= $ Dividend/market price
ie.(100*5%)/104=
4.81%
Cost of equity:
as per dividend growth model:
Ke=(D1/P0)+g
ie.((D0*(1+g))/P0)+g
14.44%
Cost of convertible bond:
is the rate which equates the after-tax cash flows on the bond & its conversion into 2 shares on maturity at the dividend growth rate
After-tax coupon interest on the bond=8%*(1-30%)=5.6%
Coupon amt.=1000*5.6%= $ 56
ie.Current MV+5 yrs.after-tax coupons+Value at maturity=0(Just like IRR calculations)
ie. -950+(56*(1-(1+r)^-5)/r)+(450*(1+3%)^5*2/(1+r)^5)=0
Solving the above, we get, the after-tax cost of debt,
rd=7.58%
Now calculating the WACC,
Type of capital Mkt. Value Propn. to total Cost Prop.*Cost
Preferred capital 15000000/100*104= 15600000 3.19% 4.81% 0.15%
Equity 10000000/10*450= 450000000 91.96% 14.44% 13.28%
Convertible debt 950/1000*25000000= 23750000 4.85% 7.58% 0.37%
Total 489350000 1 WACC= 13.80%
3.1.The firm’s weighted average cost of capital (WACC)= 13.80%
3.2 ..Dividend pay-outs certainly have their effect on share prices by signalling to the existing as well as potential investors about many parameters such as:
cash flows within the company
profitability of operations or accumulation of prior earnings
concern for shareholders and return for their money invested
sharing the efficiency of operations with the shareholders
growth prospects
They are future cash flows from holding the shares . So shareholders in need of regular income will opt for regular dividend-paying companies & hence demand for such shares increase.